# Analyzing Uncertainty Quantification in Statistical and Deep Learning Models for Probabilistic Electricity Price Forecasting

### v2

@@ 4.4 Error Metrics
with $L_{d,h}^{(1-\alpha)}$ and $U_{d,h}^{(1-\alpha)}$ denoting the lower and upper bounds of the $(1-\alpha)$ prediction interval [7]. The PICP measures the proportion of actual values that fall within the predicted interval.
@@
where $ACE_{1-\alpha}=PICP_{1-\alpha}-PINC_{1-\alpha}$ is the average calibration error, and $PINC$ is the prediction interval nominal coverage [7]. The sharpness of the predicted interval is measured using the CRPS{CRPS}, which we calculate with the pinball loss function [33]:
@@
$PS_{d,h}^{q}=\begin{cases}(1-q)(\hat{p}_{d,h}-p_{d,h}),&\text{if }p_{d,h}<\hat{p}_{d,h},\\ q(p_{d,h}-\hat{p}_{d,h}),&\text{otherwise}.\end{cases}$
The scaling factor of 2 is commonly omitted in practice [9], which we also follow here.
@@
Finally, we statistically compare model accuracies using the Diebold–Mariano (DM) test [35], with CRPS as the loss function. For the mathematical formulation of the DM test, please refer to Appendix .1.
@@ 4.6 Empirical Analysis of Diebold-Mariano Test Results
Figure 6: Heatmap of pairwise Diebold–Mariano test p-values using CRPS loss. Values below 0.05 mark significant differences in predictive accuracy.
@@
The empirical p-values are shown as a heatmap in Fig. 6, where each entry $p_{ij}$ denotes the p-value from the DM test comparing model $M_{i}$ (row model) against model $M_{j}$ (column model). In this section and the following one, we have not used five models for the Diebold-Mariano (DM) test evaluation and in the trading strategies, namely, Naive-HStrain, Ens5, MCD10, EvDNN, and EvDNN-CP, since we used only one better-performing model from each category, and the EvDNN model struggled to capture uncertainty. The diagonal elements are marked as nan since models are not tested against themselves. The p-values in the DM test show whether the two models under consideration are significantly different in terms of prediction accuracy. However, they do not provide any information about model superiority. To extract this information, one has to analyze the test statistic. The mathematical form of the DM test and the table of the test statistic are given in Appendix .1 in Table 5. Table 5 reports the corresponding DM statistics $d_{i,j}$ (based on CRPS loss), where $d_{i,j}<0$ means $M_{i}$ attains lower average loss than $M_{j}$ (row is better), $d_{i,j}>0$ means the opposite (column is better), $|d_{i,j}|$ quantifies the strength of the advantage, and values numerically near 0 indicate practical equivalence. In Fig. 6, we show the pairwise p-values (row model $M_{i}$ vs. column model $M_{j}$), with significance level $0.05$. Entries whose numeric values on the color bar are <0.05<0.05 indicate that the observed difference between the two models is statistically significant, whereas larger p-values >0.05>0.05 indicate no detectable difference.
@@
In Figure 6 we show the P-values for the DM-test computed at $\alpha=0.05$. It shows that three LEAR–type models form a statistically tied top tier: LEAR–GARCH, LEAR–QRA, and LEAR–CP are mutually indistinguishable, with near–zero differentials and non–rejections in all pairwise directions. The pair wise dm-statistic given in the Table 5 in Appendix .2 supports this claim. For further clarity on the P-values, we also give the numerical values of the P-values in Table 4 (e.g., ${DM}({\text{LEAR-GARCH},\text{LEAR-QRA}})=-0.072$ with $p=0.471$, ${DM}({\text{LEAR-CP},\text{LEAR-GARCH}})=0.049$ with $p=0.520$, and ${DM}({\text{LEAR-CP},\text{LEAR-QRA}})=-0.039$ with $p=0.484$; see Tables 54). $DM(.,.)$ represents the statistic for two models into consideration and is computed as shown in the Appdendix .1 Each of these three models, however, dominates every non-LEAR alternative by a wide and statistically decisive margin: for instance, ${DM}({\text{LEAR-CP},\text{Ens10}})=-4.921$ with $p=0.000$, ${DM}({\text{LEAR-GARCH},\text{DDNN}})=-5.763$ with $p=0.000$, and ${DM}({\text{LEAR-QRA},\text{MCD30-CP}})=-4.839$ with $p=0.000$, and analogous results hold across all such pairs. All non-LEAR models (DDNN, Ens10, MCD30) and their conformal/MC-dropout variants (DDNN-CP, Ens10-CP, MCD30-CP) are statistically indistinguishable from one another at $\alpha=0.05$, however, the directional tendencies do exist (e.g., ${DM}({\text{MCD30},\text{Ens10}})=-1.748$; $({DM}{\text{Ens10},\text{DDNN}})=-2.432$), but the associated pp–values are too large to claim superiority ($p=0.130$ and $p=0.192$, respectively). Conformal models and the MC-dropout models does not yield measurable gains (in terms of CRPS loss) over the corresponding base models: ${DM}({\text{MCD30-CP},\text{MCD30}})=0.438$ with $p=0.624$, ${DM}({\text{Ens10-CP},\text{Ens10}})=0.299$ with $p=0.578$, and ${DM}({\text{DDNN-CP},\text{DDNN}})=-0.495$ with $p=0.422$ all fail to reject. Finally, every method strictly improves upon the naive baseline (e.g., ${DM}({\text{DDNN},\text{Naive-HS}_{val}})=-6.273$, ${DM}({\text{Ens10},\text{Naive-HS}_{val}})=-6.984$, ${DM}(\text{LEAR-CP}$, $\text{Naive-HS}_{val})=-10.235$, all with $p=0.000$). In summary there is no single best model within the LEAR family, but they constitute a statistically superior frontier; (ii) non-LEAR methods are statistically tied with one another and uniformly inferior to LEAR; and (iii) naive forecasting is decisively underperforming.
@@ 4.7 Trading Strategies
Figure 7: Total profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy. For comparison, the fixed hours, optimal bids, and unlimited bids strategies are also included. Results are averaged over 10 independent runs.
@@
Figure 9: Per-transaction profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy. For comparison, the fixed hours, optimal bids, and unlimited bids strategies are also included. Results are averaged over 10 independent runs.
@@
The tables for Fig. 7, Fig. 8, and Fig. 9, along with their corresponding plots against PICP, can be found in Appendix .2.
@@ 5 Discussion and Critical Reflection
In general, it shows that there is no superior method that performs best across all error measures. However, there is a certain tendency towards LEAR-based models, which show surprisingly good performance in the presence of state-of-the-art machine learning benchmarks, despite being much simpler and requiring considerably less computational power. This indicates, at least given the setup of this case study, that the additional complexity of machine learning models is not paying off. The case study shows that the LEAR-CP model has, on average, the most reliable PIs across all models, with a MAACE of $1.305$, whereas the best machine learning model, MC Dropout, achieves a MAACE of $1.735$. At a confidence level of around $85\%$ and higher, the LEAR-QRA starts to outperform the MD model, achieving a better PICP. The evidential neural network failed to capture the uncertainty. One possible reason could be the zero-gradient problem for evidential regression models; see [36] for more information. The LEAR-GARCH also had difficulties in providing reliable PIs, as did the naive approach. The problem is that their uncertainty predictions solely depend on a historical price period.
@@
We find that improvements in reliability come at the cost of wider PIs, i.e., reduced sharpness. The naive model produced overconfident, i.e., too narrow, PIs, whereas NN-based models tend to be underconfident. However, using ensembles and MC dropout considerably improves PICP values, particularly at high confidence levels, and reduces PICP variability across runs.
@@
In the statistical models, we implemented a rolling window approach for calibrating the parameters. Although this is beneficial in many situations, it causes problems during structural breaks. Moreover, the window size needs to be chosen beforehand, and this choice is always somewhat arbitrary. Statistical models are lighter and interpretable but can be brittle under regime shifts and heavy tails if distributional assumptions are too restrictive. ML models are adaptive but need careful regularization and calibration to avoid over- or under-confidence. Rather than choosing one, the evidence suggests a hybrid path: use LEAR-type models for baseline calibration and coverage control, and combine them with deep learning models when diagnostic plots indicate systematic misspecification.
In this study, Gaussian innovations were used for comparability. Moreover, a natural extension is to test heavy-tailed and skewed families (Student-$t$, skew-$t$, generalized error, generalized hyperbolic) within LEAR-GARCH. This could retain the calibration advantages while improving tail coverage, again providing a way to strengthen statistical models without diminishing the role of ML.
@@ Acknowledgment
Part of this work is derived from the first author’s M.Sc. thesis [38] and has been substantially extended and revised for this article. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.

### v3

@@ 4.4 Error Metrics
with $L_{d,h}^{(1-\alpha)}$ and $U_{d,h}^{(1-\alpha)}$ denoting the lower and upper bounds of the $(1-\alpha)$ PI [8]. The PICP measures the proportion of actual values that fall within the predicted interval.
@@
where $ACE_{1-\alpha}=PICP_{1-\alpha}-PINC_{1-\alpha}$ is the average calibration error, and $PINC$ is the prediction interval nominal coverage [8]. The sharpness of the predicted quantiles is measured using the CRPS{CRPS}, which we calculate with the pinball loss function [8]:
@@
$PS_{d,h}^{q}=\begin{cases}(1-q)(\hat{p}_{d,h}^{q}-p_{d,h}),&\text{if }p_{d,h}<\hat{p}^{q}_{d,h},\\ q(p_{d,h}-\hat{p}_{d,h}^{q}),&\text{otherwise}\end{cases}$
where $\hat{p}^{q}_{d,h}$ denotes the forecast of the $q$th quantile of $p_{d,h}$. The scaling factor of 2 is commonly omitted in practice [10], a convention we follow here.
@@
Finally, we compare model accuracies pairwise using the Diebold-Mariano (DM) test [26], with CRPS as the loss function.
For the mathematical formulation of the DM test, please refer to Appendix A.1.
@@ 4.6 Empirical Analysis of Diebold-Mariano Test Results
Now we discuss the results of applying the Diebold-Mariano test (’DM test’), see [26] for the test. In this and in the following section, we leave out the following five models from our analysis Naive-HStrain, Ens5, MCD10, EvDNN, and EvDNN-CP, since we only want to consider well-performing models from each category. We choose a significance level of $\alpha=0.05$. We decide to not do a multiple testing correction (e.g. Bonferroni).
@@
The $p$-values in the DM test show whether the two models under consideration are significantly different in terms of prediction accuracy. However, they do not provide any information about models performance superiority. To extract this information, one has to analyze the test statistic. The mathematical form of the DM test and the table of the test statistic are given in Appendix A.1 in Table 8. Table 8 reports the corresponding DM statistics $d_{i,j}$ (based on CRPS loss), where $d_{i,j}<0$ means $M_{i}$ attains lower average loss than $M_{j}$ (row is better), $d_{i,j}>0$ means the opposite (column is better), $|d_{i,j}|$ quantifies the strength of the advantage, and values numerically near 0 indicate practical equivalence. In Fig. 6, we show the pairwise $p$-values (row model $M_{i}$ vs. column model $M_{j}$), with significance level $0.05$. Entries whose numeric values on the color bar are <0.05<0.05 indicate that the observed difference between the two models is statistically significant, whereas larger $p$-values >0.05>0.05 indicate no detectable difference. The diagonal elements are left blank since models are not tested against themselves.
Figure 6: Heatmap of pairwise Diebold–Mariano test $p$-values using CRPS loss. Values below 0.05 mark significant differences in predictive accuracy. Results are averaged over 10 independent runs, yielding 100 comparisons per model pair across runs.
@@
Each model is compared with every other model across 10 independent runs. The results show that the three LEAR–type models form a statistically tied top tier: LEAR–GARCH, LEAR–QRA, and LEAR–CP performances are mutually indistinguishable, with near-zero differentials and non-rejections in all pairwise directions. The pairwise DM statistics given in Table 8 in Appendix A.2 support this claim. For further clarity on the $p$-values, we also provide their numerical values in Table 9 in the Appendix. However, each of these three models dominates every non-LEAR alternative by a wide and statistically significant margin. All non-LEAR models (DDNN, Ens10, MCD30) and their conformal/MC Dropout variants (DDNN-CP, Ens10-CP, MCD30-CP) are statistically indistinguishable from one another at $\alpha=0.05$; however, directional tendencies do exist as MCD30 seems to be performing better, generally, although this difference is not significant. Conformal and MC Dropout models do not yield measurable gains (in terms of CRPS) over their corresponding base models. Finally, every method strictly improves upon the naive baseline.
In summary, (i) there is no single best model within the LEAR family, but together they form a statistically superior performance frontier; (ii) non-LEAR methods are statistically tied with one another and uniformly inferior to the LEAR models; and (iii) naive forecasting is decisively outperformed by all other approaches.
@@ 4.7 Trading Strategies
Figure 7: Total profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs.
@@
Figure 9: Per-transaction profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs.
@@
The tables for Fig. 7, Fig. 8, and Fig. 9, along with their corresponding plots against PICP, can be found in Appendix A.2.
@@ 5 Discussion and Critical Reflection
In general, it shows that there is no superior method that performs best across all error measures. However, there is a certain tendency towards LEAR-based models, which show surprisingly good performance in the presence of state-of-the-art machine learning benchmarks, despite being much simpler and requiring considerably less computational power. This indicates, at least given the setup of this case study, that the additional complexity of machine learning models is not paying off. The case study shows that the LEAR-CP model has, on average, the most reliable PIs across all models, with a MAACE of $1.305$, whereas the best machine learning model, MC Dropout, achieves a MAACE of $1.735$. At a confidence level of around $85\%$ and higher, the LEAR-QRA starts to outperform the MD model, achieving a better PICP. The evidential neural network failed to capture the uncertainty. One possible reason could be the zero-gradient problem for evidential regression models; see [28] for more information. The LEAR-GARCH also had difficulties in providing reliable PIs, as did the naive approach. The problem is that their uncertainty predictions solely depend on a historical price period. The superior performance of the LEAR-based models may be explained by the fact that they are refitted for every prediction. This is only possible because of the relatively low computational cost.
@@
We find that improvements in reliability come at the cost of wider PIs, i.e., reduced sharpness. The naive model produced underconfident, i.e., too wide PIs, whereas NN-based models tend to be overconfident. However, using ensembles and MC dropout considerably improves PICP values, particularly at high confidence levels, and reduces PICP variability across runs.
@@
In the statistical models, we implemented a rolling window approach for calibrating the parameters. Although this is beneficial in many situations, it causes problems during structural breaks. Moreover, the window size needs to be chosen beforehand, and this choice is always somewhat arbitrary. Statistical models are lighter and interpretable but can be brittle under regime shifts and heavy tails if distributional assumptions are too restrictive. ML models are adaptive but need careful regularization and calibration to avoid over- or underconfidence. Rather than choosing one, the evidence suggests a hybrid path: use LEAR-type models for baseline calibration and coverage control, and combine them with deep learning models when diagnostic plots indicate systematic misspecification.
In this study, Gaussian innovations were used for comparability. Moreover, a natural extension is to test heavy-tailed and skewed families (Student-$t$, skew-$t$, generalized error, generalized hyperbolic or JSU) within LEAR-GARCH. This could retain the calibration advantages while improving tail coverage, again providing a way to strengthen statistical models without diminishing the role of ML.
@@ Acknowledgment
Part of this work is derived from the first author’s M.Sc. thesis [30] and has been substantially extended and revised for this article. The authors acknowledge support by the state of Baden-Württemberg through bwHPC. The third author gratefully acknowledges that this research was (partially) funded in the course of TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (520388526) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
Appendix A
A.1 The Diebold-Mariano Test
The Diebold-Mariano (DM) test is designed as follows: Let $e_{1t}$ and $e_{2t}$ denote the forecast errors of Model 1 and Model 2, respectively, at time $t$. To quantify the loss associated with these errors, a loss function $g(\cdot)$ is employed. Common choices for $g(\cdot)$ include the squared error loss, $g(e_{t})=e_{t}^{2}$, or the absolute error loss, $g(e_{t})=|e_{t}|$. The differential loss at time $t$ is then defined as $d_{t}=g(e_{1t})-g(e_{2t})$. The null hypothesis ($H_{0}$) of the DM test posits that the two forecasts exhibit identical accuracy, implying that the expected value of the loss differential is zero:
$H_{0}:E[d_{t}]=0$
The alternative hypothesis ($H_{1}$) asserts that the forecasting accuracies are distinct:
$H_{1}:E[d_{t}]\neq 0$
The DM test statistic is constructed as:
$DM=\frac{\bar{d}}{\sqrt{\widehat{\text{Var}}(\bar{d})}}$
where $\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t}$ represents the sample mean of the loss differential over $T$ observations, and $\widehat{\text{Var}}(\bar{d})$ is a consistent estimator of the variance of $\bar{d}$. This variance estimator is crucial, as it accounts for potential serial correlation and heteroskedasticity inherent in the loss differential series $d_{t}$. Under the null hypothesis, the DM test statistic asymptotically follows a standard normal distribution, i.e., $DM\sim N(0,1)$. Effective application of the DM test for forecast accuracy assessment necessitates several key prerequisites. First, accurate computation of forecast errors for each model is required, typically defined as the difference between the observed value and the forecasted value. Second, a clearly defined and appropriate loss function must be specified to quantify the cost of forecasting errors in a manner relevant to the specific application, as the selection of the loss function directly impacts the interpretation of ”accuracy.” Third, as an asymptotic test, the DM test relies on a sufficiently large sample size ($T$) to ensure that the DM test statistic converges to its asymptotic normal distribution, thereby validating the statistical inference. Finally, while the original formulation assumed covariance stationarity of the loss differential series ($d_{t}$), modern extensions of the DM test can accommodate non-stationary series, enhancing its applicability to a broader range of time series data. The $p$-value derived from the DM test statistic serves as the basis for inference. A $p$-value below a pre-specified significance level ($\alpha$, commonly 0.05) leads to the rejection of $H_{0}$, indicating a statistically significant difference in forecasting accuracy between the two models.
A.2 Supplementary Graphics and Tables
This appendix provides additional results for Section 4. Figures 1013 show MPIW, total profit, the number of profitable days, and per-transaction profit plotted against PICP. Tables 4-9 report optimal hyperparameters, total profit, the number of profitable days, per-transaction profit, DM statistics, and DM $p$-values.
Table 4: Hyperparameters and corresponding search ranges considered during tuning for each model type.
LEAR(-GARCH/QRA) DDNN MCD(10/30) EvDNN
Learning rate - $10^{-5},10^{-1}] [$$10^{-5},10^{-1}] [$$10^{-5},10^{-1}]$
$L_{2}$ regularization parameter - $10^{-5},10^{-1}] [$$10^{-5},10^{-1}] [$$10^{-5},10^{-1}]$
Dropout rate - - $[0.01,0.9]$ -
Evidence regularization parameter - - - $10^{-4},1]$
LASSO penalty parameter $10^{-5},10^{-1}] - - -$
Figure 10: Models are evaluated using MPIW across confidence intervals ranging from 2% to 98%, and the results are plotted against PICP. The top subplot shows the average results over 10 independent runs, with standard deviations indicated by dotted lines. The bottom subplot displays the difference in performance between the DDNN and the other models, calculated using piecewise linear interpolation of the curves.
Table 5: Total profit of models is shown for confidence intervals of 50%, 60%, 70%, 80%, 90% and 98% using the quantile-based strategy. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs, with standard deviations shown in light gray. The best-performing results are highlighted in bold.
Total profit PI50% PI60% PI70% PI80% PI90% PI98%
$\text{EUR}]$
[…]
Naive-HSval 0 19280 $\pm$0 17876 $\pm$0 15866 $\pm$0 12550 $\pm$0 6669 $\pm$0 1389 $\pm$0
DDNN 0 25941 $\pm$583 25397 $\pm$900 24120 $\pm$1458 22067 $\pm$1757 17891 $\pm$2592 8491 $\pm$3551
Ens10 0 26695 $\pm$410 25907 $\pm$496 25014 $\pm$587 22786 $\pm$951 18175 $\pm$1373 5586 $\pm$1864
MCD30 0 26959 $\pm$645 26754 $\pm$554 26135 $\pm$848 23735 $\pm$1035 18346 $\pm$1644 4217 $\pm$1441
LEAR-GARCH 0 27368 $\pm$0 27652 $\pm$0 27749 $\pm$0 26004 $\pm$0 23296 $\pm$0 16485 $\pm$0
LEAR-QRA 0 28427 $\pm$0 28122 $\pm$0 28476 $\pm$0 27766 $\pm$0 25704 $\pm$0 10298 $\pm$0
DDNN-CP 0 24887 $\pm$714 23976 $\pm$873 22281 $\pm$1093 19585 $\pm$1121 14386 $\pm$1304 4699 $\pm$731
LEAR-CP 0 26757 $\pm$0 26653 $\pm$0 26388 $\pm$0 26733 $\pm$0 26224 $\pm$0 22108 $\pm$0
Ens10-CP 0 25679 $\pm$379 24964 $\pm$380 23279 $\pm$546 21065 $\pm$413 15846 $\pm$500 5349 $\pm$322
MCD30-CP 0 26111 $\pm$534 25462 $\pm$556 24299 $\pm$653 21819 $\pm$957 17212 $\pm$1315 5601 $\pm$755
Table 6: The expected number of days with profitable limit trades (satisfying Equation 2, 3, or 4) is shown for confidence intervals of 50%, 60%, 70%, 80%, 90% and 98% using the quantile-based strategy. Results are averaged over 10 independent runs, with standard deviations shown in light gray. The best-performing results are highlighted in bold.
Expected number of days with profitable limit trades PI50% 0 0 PI60% 0 0 PI70% 0 0 PI80% 0 0 PI90% 0 0 PI98% 0 0
Naive-HSval 0 65.30 $\pm$0.00 56.28 $\pm$0.00 43.99 $\pm$0.00 31.97 $\pm$0.00 16.44 $\pm$0.00 4.11 $\pm$0.00
DDNN 0 91.39 $\pm$3.09 84.87 $\pm$4.99 75.35 $\pm$6.51 64.41 $\pm$7.22 47.40 $\pm$7.77 21.01 $\pm$8.08
Ens10 0 92.50 $\pm$1.13 84.60 $\pm$1.84 75.64 $\pm$2.45 63.01 $\pm$3.18 45.37 $\pm$3.76 12.96 $\pm$4.08
MCD30 0 94.11 $\pm$1.11 89.24 $\pm$1.35 81.67 $\pm$2.36 69.07 $\pm$4.23 47.21 $\pm$4.79 10.11 $\pm$3.06
LEAR-GARCH 0 92.90 $\pm$0.00 91.26 $\pm$0.00 84.70 $\pm$0.00 76.23 $\pm$0.00 61.48 $\pm$0.00 37.43 $\pm$0.00
LEAR-QRA 0 95.36 $\pm$0.00 91.80 $\pm$0.00 88.52 $\pm$0.00 81.97 $\pm$0.00 69.67 $\pm$0.00 24.86 $\pm$0.00
DDNN-CP 0 87.02 $\pm$1.80 78.83 $\pm$2.68 68.55 $\pm$3.47 55.44 $\pm$3.57 37.79 $\pm$3.02 10.79 $\pm$1.72
LEAR-CP 0 92.90 $\pm$0.00 91.53 $\pm$0.00 89.62 $\pm$0.00 86.89 $\pm$0.00 82.51 $\pm$0.00 60.66 $\pm$0.00
Ens10-CP 0 89.34 $\pm$0.59 83.11 $\pm$0.78 71.89 $\pm$1.37 59.78 $\pm$1.11 41.12 $\pm$1.11 12.62 $\pm$0.70
MCD30-CP 0 91.01 $\pm$0.78 85.05 $\pm$1.03 76.28 $\pm$1.38 63.33 $\pm$2.84 44.86 $\pm$2.87 12.76 $\pm$1.67
Table 7: Per-transaction profit of the models is shown for confidence intervals of 50%, 60%, 70%, 80%, 90% and 98% using the quantile-based strategy. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs, with standard deviations shown in light gray. The best-performing results are highlighted in bold.
Per-transaction profit PI50% PI60% PI70% PI80% PI90% PI98%
$\text{EUR}]$
[…]
Naive-HSval 0 44.43 $\pm$0.00 46.55 $\pm$0.00 50.21 $\pm$0.00 54.56 $\pm$0.00 59.54 $\pm$0.00 46.29 $\pm$0.00
DDNN 0 40.80 $\pm$1.12 42.70 $\pm$1.48 45.50 $\pm$1.77 48.45 $\pm$1.80 52.92 $\pm$1.91 55.04 $\pm$3.58
Ens10 0 41.21 $\pm$0.45 43.42 $\pm$0.58 46.32 $\pm$0.70 50.18 $\pm$1.17 55.18 $\pm$0.68 58.71 $\pm$2.31
MCD30 0 40.81 $\pm$1.04 42.14 $\pm$1.10 44.58 $\pm$1.29 47.43 $\pm$1.30 53.62 $\pm$1.14 56.35 $\pm$4.04
LEAR-GARCH 0 43.44 $\pm$0.00 43.61 $\pm$0.00 45.79 $\pm$0.00 46.94 $\pm$0.00 51.77 $\pm$0.00 60.16 $\pm$0.00
LEAR-QRA 0 42.94 $\pm$0.00 43.94 $\pm$0.00 45.20 $\pm$0.00 47.54 $\pm$0.00 50.80 $\pm$0.00 56.58 $\pm$0.00
DDNN-CP 0 41.88 $\pm$0.99 43.83 $\pm$1.02 46.22 $\pm$0.99 49.78 $\pm$0.97 52.82 $\pm$1.75 60.29 $\pm$3.70
LEAR-CP 0 42.07 $\pm$0.00 41.78 $\pm$0.00 42.02 $\pm$0.00 42.98 $\pm$0.00 44.00 $\pm$0.00 50.02 $\pm$0.00
Ens10-CP 0 41.65 $\pm$0.43 42.73 $\pm$0.57 45.79 $\pm$0.65 49.17 $\pm$0.68 53.25 $\pm$0.98 58.02 $\pm$1.00
MCD30-CP 0 41.31 $\pm$0.98 42.48 $\pm$1.21 44.73 $\pm$0.86 47.88 $\pm$0.75 52.90 $\pm$1.10 60.45 $\pm$2.63
Table 8: Pairwise DM statistics $\widehat{DM}_{i,j}$ (row model $i$ vs. column model $j$). Negative values favor the row model, positive values favor the column model. Results are averaged over 10 independent runs, yielding 100 comparisons per model pair across runs.
Naive-HSval DDNN Ens10 MCD30 LEAR-GARCH LEAR-QRA DDNN-CP LEAR-CP Ens10-CP MCD30-CP
Naive-HSval 0 0 6.273 $\pm$0.660 6.984 $\pm$0.157 7.387 $\pm$0.356 10.619 $\pm$0.000 10.514 $\pm$0.000 6.433 $\pm$0.638 10.235 $\pm$0.000 6.898 $\pm$0.193 7.232 $\pm$0.319
DDNN 0 -6.273 $\pm$0.660 0 2.432 $\pm$2.309 3.087 $\pm$1.868 5.763 $\pm$0.712 5.781 $\pm$0.692 0.495 $\pm$2.370 5.400 $\pm$0.703 2.037 $\pm$2.179 2.650 $\pm$1.801
Ens10 0 -6.984 $\pm$0.157 -2.432 $\pm$2.309 0 1.748 $\pm$1.396 5.150 $\pm$0.231 5.184 $\pm$0.240 -1.778 $\pm$2.084 4.784 $\pm$0.225 -0.299 $\pm$1.419 1.201 $\pm$1.244
MCD30 0 -7.387 $\pm$0.356 -3.087 $\pm$1.868 -1.748 $\pm$1.396 0 4.683 $\pm$0.361 4.692 $\pm$0.388 -2.700 $\pm$1.782 4.313 $\pm$0.358 -1.856 $\pm$1.312 -0.438 $\pm$1.555
LEAR-GARCH 0 -10.619 $\pm$0.000 -5.763 $\pm$0.712 -5.150 $\pm$0.231 -4.683 $\pm$0.361 0 -0.072 $\pm$0.000 -5.731 $\pm$0.658 -0.049 $\pm$0.000 -5.264 $\pm$0.219 -4.815 $\pm$0.340
LEAR-QRA 0 -10.514 $\pm$0.000 -5.781 $\pm$0.692 -5.184 $\pm$0.240 -4.692 $\pm$0.388 0.072 $\pm$0.000 0 -5.752 $\pm$0.636 0.039 $\pm$0.000 -5.305 $\pm$0.227 -4.839 $\pm$0.377
DDNN-CP 0 -6.433 $\pm$0.638 -0.495 $\pm$2.370 1.778 $\pm$2.084 2.700 $\pm$1.782 5.731 $\pm$0.658 5.752 $\pm$0.636 0 5.374 $\pm$0.658 1.656 $\pm$2.251 2.431 $\pm$1.816
LEAR-CP 0 -10.235 $\pm$0.000 -5.400 $\pm$0.703 -4.784 $\pm$0.225 -4.313 $\pm$0.358 0.049 $\pm$0.000 -0.039 $\pm$0.000 -5.374 $\pm$0.658 0 -4.921 $\pm$0.224 -4.474 $\pm$0.340
Ens10-CP 0 -6.898 $\pm$0.193 -2.037 $\pm$2.179 0.299 $\pm$1.419 1.856 $\pm$1.312 5.264 $\pm$0.219 5.305 $\pm$0.227 -1.656 $\pm$2.251 4.921 $\pm$0.224 0 1.528 $\pm$1.336
MCD30-CP 0 -7.232 $\pm$0.319 -2.650 $\pm$1.801 -1.201 $\pm$1.244 0.438 $\pm$1.555 4.815 $\pm$0.340 4.839 $\pm$0.377 -2.431 $\pm$1.816 4.474 $\pm$0.340 -1.528 $\pm$1.336 0
Table 9: Pairwise DM p-values $p_{i,j}$ (row model $i$ vs. column model $j$). Results are averaged over 10 independent runs, yielding 100 comparisons per model pair across runs. For notational convenience, we write $p=0.000$ whenever $p<0.001$.
Naive-HSval DDNN Ens10 MCD30 LEAR-GARCH LEAR-QRA DDNN-CP LEAR-CP Ens10-CP MCD30-CP
Naive-HSval 0 0 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000 1.000 $\pm$0.000
DDNN 0 0.000 $\pm$0.000 0 0.808 $\pm$0.299 0.930 $\pm$0.181 1.000 $\pm$0.000 1.000 $\pm$0.000 0.578 $\pm$0.413 1.000 $\pm$0.000 0.776 $\pm$0.312 0.903 $\pm$0.194
Ens10 0 0.000 $\pm$0.000 0.192 $\pm$0.299 0 0.870 $\pm$0.278 1.000 $\pm$0.000 1.000 $\pm$0.000 0.243 $\pm$0.324 1.000 $\pm$0.000 0.422 $\pm$0.349 0.809 $\pm$0.277
MCD30 0 0.000 $\pm$0.000 0.070 $\pm$0.181 0.130 $\pm$0.278 0 1.000 $\pm$0.000 1.000 $\pm$0.000 0.090 $\pm$0.203 1.000 $\pm$0.000 0.118 $\pm$0.268 0.376 $\pm$0.289
LEAR-GARCH 0 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0 0.471 $\pm$0.000 0.000 $\pm$0.000 0.480 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000
LEAR-QRA 0 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0.529 $\pm$0.000 0 0.000 $\pm$0.000 0.516 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000
DDNN-CP 0 0.000 $\pm$0.000 0.422 $\pm$0.413 0.757 $\pm$0.324 0.910 $\pm$0.203 1.000 $\pm$0.000 1.000 $\pm$0.000 0 1.000 $\pm$0.000 0.722 $\pm$0.348 0.882 $\pm$0.220
LEAR-CP 0 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0.000 $\pm$0.000 0.520 $\pm$0.000 0.484 $\pm$0.000 0.000 $\pm$0.000 0 0.000 $\pm$0.000 0.000 $\pm$0.000
Ens10-CP 0 0.000 $\pm$0.000 0.224 $\pm$0.312 0.578 $\pm$0.349 0.882 $\pm$0.268 1.000 $\pm$0.000 1.000 $\pm$0.000 0.278 $\pm$0.348 1.000 $\pm$0.000 0 0.846 $\pm$0.274
MCD30-CP 0 0.000 $\pm$0.000 0.097 $\pm$0.194 0.191 $\pm$0.277 0.624 $\pm$0.289 1.000 $\pm$0.000 1.000 $\pm$0.000 0.118 $\pm$0.220 1.000 $\pm$0.000 0.154 $\pm$0.274 0
Figure 11: Total profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy, and the results are plotted against PICP. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs.
Figure 12: The expected number of days with profitable limit trades (satisfying Equation 2, 3, or 4) is shown for each model across confidence intervals ranging from 2% to 98% using the quantile-based strategy and the results are plotted against PICP. Results are averaged over 10 independent runs.
Figure 13: Per-transaction profit of models is shown for confidence intervals ranging from 2% to 98% using the quantile-based strategy, and the results are plotted against PICP. For comparison, the optimal bids strategy is also included. Results are averaged over 10 independent runs.